# Multimodal Trajectory Representation Learning for Travel Time Estimation

### v1

@@ 4.6. Self-supervised Training
(18) $\mathcal{L}_{CL}=\frac{1}{2}\left[\mathcal{L}_{\tilde{G}\rightarrow R}+\mathcal{L}_{R\rightarrow\tilde{G}}\right$],
@@
(19) $\mathcal{L}_{\tilde{G}\rightarrow R}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\tilde{G}_{i}\cdot R_{i}/\tau)}{\sum_{j=1}^{N}\exp(\tilde{G}_{i}\cdot R_{j}/\tau)},$
(20) $\mathcal{L}_{R\rightarrow\tilde{G}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(R_{i}\cdot\tilde{G}_{i}/\tau)}{\sum_{j=1}^{N}\exp(R_{i}\cdot\tilde{G}_{j}/\tau)}.$
@@ 5.1. Experiment Settings
Due to space limitations, we have placed RQ4 in the Appendices.
@@
Implementation Details. The proposed MDTI model integrates a GPS encoder, a grid encoder, a road encoder, a dynamic trajectory alignment module, and a cross-modal interaction module, trained in an end-to-end manner. Training is performed with a batch size of 32 for 30 epochs. We adopt Adam (Kingma, 2014) as the optimizer with an initial learning rate of $2\times 10^{-4}$, cosine annealing schedule, 10 warm-up epochs, a minimum learning rate of $1\times 10^{-6}$, and weight decay of $1\times 10^{-4}$. Both hidden and output embedding dimensions are fixed at 256 with a dropout rate of 0.1. For GPS trajectories, each sequence is segmented into patterns and compared with historical keys, and the most relevant ones are transformed into natural language prompts. These prompts are encoded by a pre-trained GPT-2 model, where the hidden state of the last token is taken as the trajectory embedding. Grid and road representations are generated by Transformer and GAT layers, while the cross-modal module aligns multi-source features. The objective combines multiple loss terms with weights of relation=0.5, step=1.0, rel=1.0, and cos=1.0.
Evaluation Metrics. We adopt three widely-used regression metrics: Mean Absolute Error (MAE) for measuring average prediction error magnitude, Mean Absolute Percentage Error (MAPE) for assessing relative error percentage. These metrics collectively provide a comprehensive assessment of prediction accuracy and model robustness.
Baselines. We compare the proposed MDTI model with eight baseline methods, which can be categorized into four groups. All methods are trained using the same number of trajectories. The four groups include: one method based on GPS trajectories (e.g., Traj2vec (Yao et al., 2017b)); one method based on grid trajectories (e.g., TrajCL (Chang et al., 2023)); four methods based on road trajectories (e.g., JCLRNT (Mao et al., 2022), START (Jiang et al., 2023), JGRM (Ma et al., 2024), GREEN (Zhou et al., 2025)); and two methods related to travel time estimation (e.g., DutyTTE (Mao et al., 2025) and MulT-TTE (Liao et al., 2024)). Further details about the baselines are provided in Appendices 7.1.
@@ 5.2. Performance Comparison (RQ1)
Table 1 presents the overall performance of our model and baseline methods across the three city datasets. Several key observations emerge: (1) Our method consistently achieves the best MAE on all datasets and the lowest MAPE on Chengdu and Xi’an. This confirms that the proposed cross-modal interaction module and modality-specific encoders within a unified feature space effectively enhance cross-domain alignment and semantic representation. In addition, the dynamic trajectory feature extraction mechanism adaptively balances information density across trajectories of varying lengths, mitigating local embedding bias. (2) Among single-modality trajectory modeling methods, Traj2vec—which directly encodes raw GPS point sequences—performs the worst, as RNN-based sequence modeling amplifies local noise. In contrast, TrajCL, which relies solely on grid representations, suffers from spatial discretization errors and the loss of fine-grained positional information. Road-network-based methods generally outperform both, indicating that the topological connectivity of road segments and intersections provides strong structural constraints, while road semantics and traffic rules capture more stable spatiotemporal patterns. (3) On the Porto dataset, our method performs slightly worse than MulT-TTE in terms of MAPE, likely due to MAPE’s sensitivity to relative errors. Nevertheless, our approach achieves the lowest MAE, demonstrating superior robustness to outliers and measurement noise. DutyTTE shows the weakest overall performance, as it lacks multimodal alignment and focuses primarily on global path consistency and interval coverage. (4) Overall, cross-modal trajectory representation learning methods outperform single-modality ones, confirming that multimodal information effectively compensates for modality-specific limitations. Furthermore, our dynamic fusion mechanism adaptively regulates information density according to trajectory length, preventing feature over-concentration in local segments. This design yields consistently stable performance across both short and long trajectories while reducing the impact of extreme values.
@@
GPS-based trajectory representation learning
Traj2vec 4.027 0.534 1.651 0.352 2.481 0.267
Grid-based trajectory representation learning
TrajCL 2.085 0.236 1.632 0.321 1.609 0.536
Travel Time Estimation Models
DutyTTE 3.780 0.303 4.131 0.302 4.525 0.316
MulT-TTE 1.524 0.147 1.231 0.172 1.461 0.234
Road-based trajectory representation learning
START 1.541 0.173 1.764 0.282 1.359 0.218
JCLRNT 3.065 0.327 1.638 0.272 1.352 0.241
JGRM 2.993 0.375 1.457 0.333 1.374 0.360
GREEN 1.510 0.165 1.241 0.184 1.322 0.185
MDTI 1.419 0.152 1.169 0.170 1.143 0.183
@@ 5.3. Component Contribution (RQ2)
Ablation Settings. To assess the contribution of each MDTI module, we conduct four ablations:
@@
  • w/o Dynamic Fusion This variant removes the prompt-based Dynamic Fusion mechanism.
The outcomes of the ablation study conducted on the Porto dataset, as reported in Table 2, quantitatively validate the contribution of each component within the MDTI framework. Additional analyses and extended ablation results are provided in the Appendices 7.2.
@@
5.4. Trajectory Length Impact (RQ3)
@@
6. CONCLUSION

### v2

@@ 3.6. Self-supervised Training
(19) $\mathcal{L}_{CL}=\frac{1}{2}\left[\mathcal{L}_{\tilde{G}\rightarrow\mathbf{R}}+\mathcal{L}_{\mathbf{R}\rightarrow\tilde{G}}\right$],
@@
(20) $\mathcal{L}_{\tilde{G}\rightarrow\mathbf{R}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\tilde{G}_{i}\cdot\mathbf{R}_{i}/\tau)}{\sum_{j=1}^{N}\exp(\tilde{G}_{i}\cdot\mathbf{R}_{j}/\tau)},$
(21) $\mathcal{L}_{\mathbf{R}\rightarrow\tilde{G}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\mathbf{R}_{i}\cdot\tilde{G}_{i}/\tau)}{\sum_{j=1}^{N}\exp(\mathbf{R}_{i}\cdot\tilde{G}_{j}/\tau)}.$
@@ 4.1. Experiment Settings
Due to space limitations, we have placed RQ3 and RQ4 in the Appendix.
@@
Implementation Details. Training is performed with a batch size of 32 for 30 epochs. We adopt Adam (Kingma, 2014) as the optimizer with an initial learning rate of $2\times 10^{-4}$, cosine annealing schedule, 10 warm-up epochs, a minimum learning rate of $1\times 10^{-6}$, and weight decay of $1\times 10^{-4}$. Both hidden and output embedding dimensions are fixed at 256 with a dropout rate of 0.1. For GPS trajectories, each sequence is segmented into patterns and compared with historical keys, and the most relevant ones are transformed into natural language prompts. These prompts are encoded by a pre-trained GPT-2 model, where the hidden state of the last token is taken as the trajectory embedding. Grid and road representations are generated by Transformer and GAT layers, while the cross-modal module aligns multi-source features. The objective combines multiple loss terms with weights of relation=0.5, step=1.0, rel=1.0, and cos=1.0.
Evaluation Metrics. We adopt two widely-used regression metrics: Mean Absolute Error (MAE) for measuring average prediction error magnitude, Mean Absolute Percentage Error (MAPE) for assessing relative error percentage. These metrics collectively provide a comprehensive assessment of prediction accuracy and model robustness.
Baselines. We compare the proposed MDTI model with eight baseline methods, which can be categorized into four groups. All methods are trained using the same number of trajectories. The four groups include: one method based on GPS trajectories (e.g., Traj2vec (Yao et al., 2017a)); one method based on grid trajectories (e.g., TrajCL (Chang et al., 2023)); four methods based on road trajectories (e.g., JCLRNT (Mao et al., 2022), START (Jiang et al., 2023), JGRM (Ma et al., 2024), GREEN (Zhou et al., 2025)); and two methods related to travel time estimation (e.g., DutyTTE (Mao et al., 2025) and MulT-TTE (Liao et al., 2024)). Further details about the baselines are provided in Appendix A.2.
@@ 4.2. Performance Comparison (RQ1)
Table 1 presents the overall performance of our model and baseline methods across the three city datasets. Several key observations emerge: (1) Our method consistently achieves the best MAE on all datasets and the lowest MAPE on Chengdu and Xi’an. This confirms that the proposed cross-modal interaction module and modality-specific encoders within a unified feature space effectively enhance cross-domain alignment and semantic representation. In addition, the dynamic trajectory feature extraction mechanism adaptively balances information density across trajectories of varying lengths, mitigating local embedding bias. (2) Among single-modality trajectory modeling methods, Traj2vec–which directly encodes raw GPS point sequences–performs the worst, as RNN-based sequence modeling amplifies local noise. In contrast, TrajCL, which relies solely on grid representations, suffers from spatial discretization errors and the loss of fine-grained positional information. Road-network-based methods generally outperform both, indicating that the topological connectivity of road segments and intersections provides strong structural constraints, while road semantics and traffic rules capture more stable spatiotemporal patterns. (3) On the Porto dataset, our method performs slightly worse than MulT-TTE in terms of MAPE, likely due to MAPE’s sensitivity to relative errors. Nevertheless, our approach achieves the lowest MAE, demonstrating superior robustness to outliers and measurement noise. DutyTTE shows the weakest overall performance, as it lacks multimodal alignment and focuses primarily on global path consistency and interval coverage. (4) The cross-modal trajectory representation learning methods outperform single-modality ones, confirming that multimodal information effectively compensates for modality-specific limitations. Our dynamic fusion mechanism adaptively regulates information density according to trajectory length, preventing feature over-concentration in local segments. This design yields consistently stable performance across both short and long trajectories while reducing the impact of extreme values.
@@ 4.3. Component Contribution (RQ2)
Ablation Settings. To assess the contribution of each MDTI module, we conduct six ablations:
@@
  • w/o Dynamic Fusion. This variant removes the prompt-based Dynamic Fusion mechanism.
  • w/o CL Loss. Removes the contrastive loss.
  • w/o MLM Loss. Removes the MLM loss.
The outcomes of the ablation study conducted on the Porto dataset, as reported in Table 2, quantitatively validate the contribution of each component within the MDTI framework.
@@
Effectiveness of Pre-training Objectives. Ablating the two self-supervised pre-training tasks also confirms their essential roles. Removing the MLM loss leads to a more pronounced degradation than removing the CL loss, suggesting that the MLM task—which reconstructs masked road segments based on contextual and grid information—plays a key role in modeling intra-modal dependencies and cross-modal interactions. The relatively smaller impact of removing the CL loss indicates that while inter-modal contrastive alignment enhances representation consistency, the model can still acquire meaningful embeddings through the MLM objective and supervised training. Nevertheless, the best results are achieved when both objectives are jointly optimized, demonstrating their complementary effects in improving representation quality and downstream performance.
@@
5. Related Work
5.1. Multimodal Trajectory Representation Learning
TRL encodes sequences of states, actions, or spatiotemporal points into compact vectors that preserve temporal structures and semantic information. Early work introduced Traj2vec (Yao et al., 2017a), which directly modeled raw GPS trajectories as point sequences to derive compact representations for downstream tasks. However, due to inherent issues in GPS data—such as positioning drift, U-turns, and redundant sampling—these approaches often produced unstable and poorly transferable representations (Li et al., 2008).
To overcome these limitations, grid-based trajectory modeling methods have been proposed (Chang et al., 2023; Li et al., 2018). By discretizing the spatial domain into grid cells and constructing graph-based structures, these approaches enhance the capacity for semantic and structural trajectory modeling. For instance, TrajCL (Chang et al., 2023) adopts contrastive learning to model spatial correlations and structural relationships among trajectories jointly.
In parallel, road-based trajectory modeling has emerged as another major direction. Using map-matching techniques (Lou et al., 2009), trajectories are aligned with road networks, which introduces topological constraints and enables richer representations of travel patterns. JCLRNT (Mao et al., 2022) generates contrastive samples between roads by leveraging contextual neighbors and constructing trajectory-level contrasts through detour path replacement. START (Jiang et al., 2023) incorporates temporal regularities and travel semantics into a self-supervised framework, improving both representation quality and generalization. JGRM (Ma et al., 2024) further strengthens spatiotemporal road features by mapping raw GPS points to corresponding road segments(Kanoulas et al., 2006; Yuan et al., 2012; Zhu et al., 2025). Based on this trend, GREEN (Zhou et al., 2025) integrates grid-based and road-based representations to exploit their complementarity.
Building on these works, we unify GPS-based, grid-based, and road-based information in a multimodal framework to mitigate single-modality limitations for TTE.
5.2. Travel Time Estimation
Travel time estimation refers to predicting the time required to traverse a route or road segment from origin to destination. Existing studies can be broadly categorized into path-based methods and origin–destination (OD)-based approaches.
Path-based methods (Ta et al., 2022; Li et al., 2023) use complete trajectories or trajectory segments as input, enabling the modeling of vehicle dynamics within the road network. These approaches typically leverage road segment features, trajectory sequences, or path structures to capture spatiotemporal dependencies. For instance, DeepTTE (Wang et al., 2018a) integrates graph convolutional networks with LSTM (Yao et al., 2017b) to model the spatiotemporal dependencies of GPS sequences. HierETA (Chen et al., 2022) introduces a multi-view hierarchical choice model for fine-grained and interpretable TTE. MulT-TTE (Liao et al., 2024) jointly models trajectory, attribute, and semantic sequences, and further applies self-supervised learning to improve generalization, yielding significant gains in prediction accuracy.
OD-based methods (Mao et al., 2025; Liu et al., 2025), by contrast, rely only on OD pairs and departure times, avoiding the need for full trajectory inputs. More recently, DutyTTE (Mao et al., 2025) combines reinforcement learning with a mixture-of-experts framework to optimize path prediction and capture segment-level uncertainty, enabling confidence interval estimation.
Our work is more closely aligned with path-based approaches, as it exploits the spatiotemporal information embedded in complete trajectories to achieve more accurate and robust TTE.
6. Conclusion
7. Acknowledgment
This work was supported in part by the Zhejiang Province Basic Public Welfare Research Project under Grant LTGG24F020009, in part by the National Natural Science Foundation of China under Grant 62476247, in part by the ”Pioneer” and ”Leading Goose” R&D Program of Zhejiang under Grant 2024C01214, in part by the Zhejiang Provincial Natural Science Foundation under Grant LR21F020003, and in part by the Zhejiang Provincial Natural Science Foundation of China under Grant No.LQN26F030027.
Appendix A Appendix
A.1. Notation
We summarize and list all notations used throughout this paper in the following Table  LABEL:notation.
Table 3. Notation definitions used in this paper.
Notation Description $\mathcal{T}^{\text{gps}}$ an ordered sequence of raw GPS points $\mathcal{T}^{\text{grid}}$ a grid trajectory $\mathcal{G}$ a road network $\mathcal{V}$ the set of vertices $\mathcal{E}$ the set of directed edges $\mathcal{T}^{\text{road}}$ a road trajectory $A$ a first-order adjacency matrix $T$ the number of temporal segments $S_{m}$ a set of sub-trajectories $D$ the embedding dimension of the Transformer input $\bar{K}$ the number of predefined reference patterns $z^{\prime}$ projected GPS embedding sequence $\mathbf{R}$ a road sequence $G$ a grid embedding sequence $\tilde{G}$ a fused grid representation
A.2. Baseline
The baseline methods employed in this paper are detailed as follows: 1. GPS-based trajectory representation learning.
  • Traj2vec (Yao et al., 2017a): An RNN-based seq2seq model converts GPS trajectory into a feature sequence to learn trajectory representations.
2. Grid-based trajectory representation learning.
  • TrajCL (Chang et al., 2023): The state-of-the-art grid-based method proposes a set of trajectory augmentations on grid trajectory in free space and dual-feature self-attention to learn grid trajectory representations using contrastive learning with the Transformer.
3. Road-based trajectory representation learning.
  • JCLRNT (Mao et al., 2022): A jointly contrastive learning framework that performs within-road, within-trajectory, and cross-scale road–trajectory contrast, maximizing mutual information to obtain robust representations.
  • START (Jiang et al., 2023): A self-supervised trajectory representation model that incorporates temporal regularities and travel semantics using a graph attention network and a time-aware trajectory encoder, trained with span-masked recovery and contrastive learning tasks.
  • JGRM (Ma et al., 2024): A joint GPS–route modeling approach with dual encoders and a cross-modal interaction module, trained in a self-supervised manner via masked language modeling (MLM) and cross-modal matching (CMM) objectives.
  • GREEN (Zhou et al., 2025): A multimodal trajectory representation method that jointly leverages grid-based and road-based trajectories, aligned and fused via contrastive learning and masked reconstruction objectives.
4. Travel Time Estimation Models.
  • DutyTTE (Mao et al., 2025): A reinforcement learning-based framework that enhances the alignment between predicted and actual travel routes, and employs a mixture-of-experts mechanism to quantify uncertainty across different road segments.
  • MulT-TTE (Liao et al., 2024): A multi-faceted model that decomposes routes into trajectory, attribute, and semantic sequences, combines sequential learning with Transformer encoders, and incorporates self-supervised masking and multi-task optimization.
Table 4. The transferability accuracy of MDTI across three cities.
[…]
Porto $\rightarrow$ Chengdu
   w/ Transfer 1.157 1.634 0.193
   w/o Transfer 1.143 1.508 0.183
Chengdu $\rightarrow$ Xi’an
   w/ Transfer 1.153 1.568 0.173
   w/o Transfer 1.169 1.614 0.170
A.3. Trajectory Length Impact (RQ3)
@@
A.4. Transferability Study (RQ4)
In Table 4, we evaluate the transferability of our pre-trained model across distinct urban scenarios through a series of experiments. In particular, we fine-tuned the model on the TTE task to adapt it to different city environments, while employing the transferred prompt files during task training.
The results demonstrate that transferring from Porto to Chengdu did not yield additional improvements; instead, the performance slightly declined. Nevertheless, the outcomes remain competitive compared with other baselines. This phenomenon may be attributed to the substantial differences in traffic patterns between the two cities, which limit the effectiveness of transfer learning. Conversely, transferring from Chengdu to Xi’an led to overall performance gains, suggesting that the two cities share more similar traffic characteristics, thereby enabling transfer learning to exert a positive effect.
A.5. Implementation Details
As a supplement to the implementation details in the main paper, we provide a detailed explanation of Algorithm 1, which describes the multimodal pretraining process of MDTI.
In the initialization stage, we set the model parameters $\Theta$ (line 1) and prepare the road network graph $\mathcal{G}_{\text{road}}$, the grid graph $\mathcal{G}_{\text{grid}}=(X,A)$, and the GPS trajectories $\{\tau_{i}\}$. Next, in Stage 1 , for each training iteration we sample a mini-batch of trajectories and, for each trajectory $\tau_{i}$ (lines 2-3), construct its road segment sequence on $\mathcal{G}_{\text{road}}$. We then update the road node features with a GAT module and feed the resulting road sequence into a Transformer to obtain the road sequence representation $R_{i}$ (lines 4-8). Then, in Stage 2 , for each trajectory we extract local grid subgraphs along its path on $\mathcal{G}_{\text{grid}}$, apply a two-layer GAT on $(X,A)$ to obtain grid node embeddings $H$, and project them to form the grid sequence embeddings $G_{i}$ (lines 9-13). In Stage 3, we further process each GPS trajectory $\tau_{i}$ by segmenting it into sub-trajectories and computing motion patterns. These patterns are organized into structured prompts $\mathrm{Prompt}_{i,t}$, which are fed into a language model $\mathrm{LM}(\cdot)$ to obtain GPS token embeddings $Z_{i}=\{z_{i,t}\}_{t=1}^{T}$ (lines 14-18). Next, in Stage 4, we align $G_{i}$ and $Z_{i}$ into the same latent space and use a residual fusion module to obtain a fused grid representation $\tilde{G}_{i}$. We then apply cross-attention between the road sequence representation $R_{i}$ and $\tilde{G}_{i}$ to produce the fused trajectory representation $\hat{R}_{i}$ (lines 19-23). In Stage 5, we compute a contrastive loss $\mathcal{L}_{\mathrm{CL}}$ between the fused grid representations and the road sequence representations (line 24), encouraging trajectories from the same instance to be close while pushing apart those from different instances. Finally, in Stage 6 , we randomly mask a subset of road tokens in $\{R_{i}\}$, predict the masked tokens with the road encoder, and compute the MLM loss $\mathcal{L}_{\mathrm{MLM}}$. The total loss $\mathcal{L}=\mathcal{L}_{\mathrm{CL}}+\mathcal{L}_{\mathrm{MLM}}$ is then used to update the model parameters $\Theta$ via gradient descent (lines 25-28), and this process is repeated until convergence.
A.6. Case Study
We randomly selected ride-hailing trips in Porto as our testbed. We simulated travel-time estimation using different approaches and compared the estimated durations with the ground-truth travel times. As shown in Figure 4, we visualize a local region of Porto using three types of base maps: grids, GPS traces, and the road network. We use color-coded arrows to indicate the predictions produced by different models.
Most baselines tend to follow routes dictated by their input trajectory representation when estimating travel time. For example, the GPS-based Traj2vec plans routes according to GPS points, whereas the grid-based TrajCL infers routes based on grid transitions. In contrast, our method jointly leverages multiple trajectory modalities to identify the most plausible route, and its travel-time estimates are the closest to the ground truth. These results demonstrate the effectiveness of the proposed approach.
Figure 4. A case study of travel time estimation.
Algorithm 1 Multimodal Pretraining Process
0: Road network graph $\mathcal{G}_{\text{road}}$, grid graph $\mathcal{G}_{\text{grid}}=(X,A)$, GPS trajectories $\{\tau_{i}\}$, prototype pattern set $\mathcal{P}$, batch size $N_{b}$, learning rate $\eta$.
0: all trained model parameters $\boldsymbol{\Theta}$.
1: Initialize parameters $\boldsymbol{\Theta}$
2:for Each training iteration do
3:  Sample a mini-batch of trajectories $\{\tau_{i}\}_{i=1}^{N_{b}}$ Stage 1: Road Trajectory Encoding:
4:  for Each trajectory $\tau_{i}$ do
5:   Construct road segment sequence $\boldsymbol{r}_{i}$ from $\tau_{i}$ on $\mathcal{G}_{\text{road}}$
6:   Update road node features by GAT on $\mathcal{G}_{\text{road}}$
7:   Encode $\boldsymbol{r}_{i}$ with Transformer to obtain road sequence representation $\boldsymbol{R}_{i}$
8:  end forStage 2: Grid Trajectory Encoding:
9:  for each trajectory $\tau_{i}$ do
10:   Extract local grid subgraphs along $\tau_{i}$ and obtain grid node features $X$
11:   Apply a two-layer GAT on $(X,A)$ to compute grid node embeddings $\boldsymbol{H}$
12:   Project $\boldsymbol{H}$ into grid sequence embeddings $\boldsymbol{G}_{i}$
13:  end forStage 3: GPS Trajectory Encoding:
14:  for Each trajectory $\tau_{i}$ do
15:   Segment $\tau_{i}$ into sub-trajectories and compute motion patterns
16:   Form structured prompts $\text{Prompt}_{i,t}$ and feed them into language model $\mathrm{LM}(\cdot)$
17:   Obtain GPS token embeddings $\boldsymbol{Z}_{i}=\{\boldsymbol{z}_{i,t}\}_{t=1}^{T}$
18:  end forStage 4: Embedding Fusion:
19:  for Each trajectory $\tau_{i}$ do
20:   Align $\boldsymbol{G}_{i}$ and $\boldsymbol{Z}_{i}$ into the same latent space
21:   Fuse them via residual fusion to get fused grid representation $\tilde{\boldsymbol{G}}_{i}$
22:   Apply cross-attention between road sequence $\boldsymbol{R}_{i}$ and $\tilde{\boldsymbol{G}}_{i}$ to obtain fused trajectory representation $\hat{\boldsymbol{R}}_{i}$
23:  end forStage 5: Contrastive:
24:  Compute contrastive loss $\mathcal{L}_{\text{CL}}$ between $\{\tilde{\boldsymbol{G}}_{i}\}$ and $\{\boldsymbol{R}_{i}\}$ Stage 6: Masked Language Modeling:
25:  Randomly mask a subset of road tokens in $\{\boldsymbol{R}_{i}\}$
26:  Predict the masked tokens with the road encoder and compute MLM loss $\mathcal{L}_{\text{MLM}}$
27:  Compute total loss $\mathcal{L}=\mathcal{L}_{\text{CL}}+\mathcal{L}_{\text{MLM}}$
28:  Update parameters $\boldsymbol{\Theta}\leftarrow\boldsymbol{\Theta}-\eta\nabla_{\boldsymbol{\Theta}}\mathcal{L}$
29:end for
30:return all trained parameters $\boldsymbol{\Theta}$